# New Permutation Decomposition Techniques For Efficient Homomorphic Permutation

### v5

@@ 3.2. Concrete Design
Based on the observation above, we design an algorithm searching a depth-1 ideal decomposition $U=U_{L}U_{R}$ for arbitrary $n\times n$ permutation $U$. The algorithm is composed of three steps.
Step 1. (Decompose Entries) This step decomposes each non-zero entry in $U^{\prime}$ to the $r_{c}$-th, $-r_{c}$-th, or 00-th diagonals of $U_{R}$ where $r_{c}$ corresponds to $r_{c}^{(1)}$ in Definition 2). First, we create a set of key-value tables representing the non-zero diagonals of $U_{R}$ :
@@
where $M_{i}$ represents the $i$-th non-zero diagonal for $i\in\{\pm r_{c},0\}$. For each table, its keys indicate the row coordinates of entries on it, and its values denote the diagonal coordinates of entries in $U$ determined by the keys. Next, for each non-zero entry $U(k,l)_{r}$ in $U^{\prime}$:
@@
    If $-r_{c}<k<r_{c}$, add $\langle\texttt{ConvertR}(k,l,0),k\rangle$ to $M_{0}$;
@@
According to Theorem 1 and Definition 2, for the diagonal distribution of $U_{L}$ to converge to $[-r+r_{c},r-r_{c}]$ with $r_{c}=a\cdot\lceil\frac{r/a}{2}\rceil$, non-zero entries on the $k$-th diagonal of $U$ with $|k|>r_{c}$ can only be decomposed to the $\texttt{sign}(k)\cdot r_{c}$-th diagonal of $U_{R}$. Step 1 employs the first two conditional statements to receive these entries for decomposition. If conflicts arise between the $r_{c}$-th and $-r_{c}$-th diagonals during the decomposition of these entries, it indicates that no ideal decomposition exists for $U$. Such cases are captured by the fourth conditional statement.
The remaining non-zero entries of $U$ are distributed on diagonals whose indices’ absolute value do not exceed $r_{c}$, and these entries can either be decomposed to the 0-th diagonal or the nearest diagonal among the $r_{c}$-th and $-r_{c}$-th diagonals according to Theorem 1. The strategy in Step 1 is to decompose all of them to the 0-th diagonal.
After Step 1, $U_{R}$ has non-zero entries only on the 0-th and ±$\pm r_{c}$-th diagonals, with potential conflicts occurring only between the 0-th and $r_{c}$-th diagonals, as well as between the 0-th and $-r_{c}$-th diagonals.
@@
Assigned with some row index $a$, the function CheckRow first checks if any conflict exists in row $a$. If there is a conflict, it attempts to resolve the conflict; otherwise, it proceeds to resolve the next conflict in $Q$. From the previous step, it is known that the conflicts occur between the 0-th diagonal and the ±$\pm r_{c}$-th diagonals. Therefore, for any conflicting row $a$, CheckRow can only resolve it by moving the entry $U_{R}(0,a)_{r}$ to the $r_{c}$-th (or $-r_{c}$-th) diagonal. The condition for successfully moving $U_{R}(0,a)_{r}$ is that the target row $\hat{a}$ must not contain any entries on the $r_{c}$-th (or $-r_{c}$-th) diagonal. If such an entry exists (with a corresponding key-value pair $\langle\hat{a},v\rangle$ in $M_{k}$ for $k=r_{c}$ or $-r_{c}$), two possible cases arise:
@@
    If $|v|>r_{c}$, then $U_{R}(0,a)_{r}$ conflicts with $U_{R}(k,\hat{a})_{r}$ after the movement, but $U_{R}(k,\hat{a})_{r}$ is not movable.
@@
    Otherwise, $U_{R}(k,\hat{a})_{r}$ must have originally been moved from the 00-th diagonal to row $\hat{a}$. If continuing to move this entry results in an ideal decomposition, by Theorem 2, there must be an alternative order of entry movements that avoids moving any entry more than once and still reaches the same solution.
Therefore, $U_{R}(0,a)_{r}$ should not be moved to row $\hat{a}$ in either of these cases. If the movement is successful, CheckRow invokes itself to resolve any potential conflict between the $r_{c}$-th (or $-r_{c}$-th) diagonal and the 00-th diagonal for row $\hat{a}$. If the invocation returns 00 for irresolvable conflicts, then CheckRow has to undo the movement it employs and return 00 to the previous recursive call. If the invocation determines no conflict exists in row $\hat{a}$, it proceeds to resolve the next conflict recorded in $Q$. In this way, CheckRow operates as a depth-first search, exhaustively checking all entry movement combinations until it identifies the first conflict-free $U_{R}$.
Theorem 2.
For any depth-1 ideal decomposition $U=U_{L}^{\prime}U_{R}^{\prime}$ of $U$, there exists an order of entry movements where each entry in $U_{R}$ is moved at most once to obtain $U_{R}^{\prime}$ after Step 1 of the search algorithm.
Proof.
For $0\leq i<n$, move the non-zero entry in the $i$-th column of $U_{R}$ to the position of the non-zero entry in the $i$-th column of $U_{R}^{\prime}$. Clearly, each entry is moved only once. ∎
@@
The correctness of the search algorithm that it detects whether the given permutation $U$ has any ideal depth-1 decomposition is demonstrated throughout the description above. In the next section, the search algorithm is applied to specific permutations in homomorphic matrix operations, where we provide a visualizable demonstration of this algorithm.
@@ 4.1. Homomorphic Matrix Transposition
Let $ct(\mathbf{a})$ denote a ciphertext of an $n$-dimensional vector $\mathbf{a}$, where $\mathbf{a}$ is the row-ordering encoding of a $d\times d$ matrix $A$ ($d^{2}\leq n$). To obtain the ciphertext of the row-ordering vector of $A^{T}$, a specific homomorphic permutation (denoted as $n\times n$ matrix $U^{t}$) is applied to $ct(\mathbf{a})$. $U^{t}$ possesses $12d-$ 1 nonzero diagonals defined as:
$U^{t}\Big{(}i(d-1),j(d+1)+ik\Big{)}_{r}=\begin{cases}1,&i\in[0,d),k=1;\\ 1,&i\in(-d,0),k=-d;\\ 0,&\text{otherwise}.\\ \end{cases}$
where $j\in[0,d-|i|)$. Currently, the BSGS-based HLT mentioned previously yields the best performance for $U^{t}$ with an overhead of approximately $d_{1}+2d_{2}$ rotations and keys, where $d_{1}d_{2}=d$. As noted previously, using the Benes network for decomposition would yield a rotation complexity no less than $O(4\log{d^{2}})$, which is inefficient within the range $0<d<2^{8}$ for practical HE parameter settings (Bossuat et al., 2024), not to mention the required $2\log{d^{2}}$ multiplicative depth.
4.1.1. Case with Dimension a Power of $2$
For $d$ a power of $2$, $U^{t}$ can be precisely decomposed to $U^{t}_{L}U^{t}_{R}$ when we apply the depth-$1$ decomposition searching algorithm proposed in Section 3. Figure 2 illustrates this process of a $16\times 16$ $U^{t}$.
The $r_{c}$ is set to $(d-1)\cdot\lceil\frac{d-1}{2}\rceil=6$ as the indices of non-zero diagonals in $U^{t}$ have a common difference of $d-1=3$. Step 1 of the algorithm is presented in Figure 2b. We can observe that $U^{t}$’s entries in diagonals $k\geq 6$ and $k\leq-6$ are decomposed to the $6$-th and $-6$-th diagonal of $U_{R}^{t}$ respectively, while those in diagonal $-6<k<6$ are decomposed to the 00-th diagonal. This introduces conflicts in rows $6$ and $9$. Figure 2c shows how Step 2 resolves the conflicts, illustrated by the dash paths. For example, let CheckRow begin with row $9$: CheckRow first moves the conflicting entry $U_{R}(0,9)$ to the $6$-th diagonal as $M_{0}[9]>0$, and this corresponds to a destination of row $3$. CheckRow then invokes itself to check row $3$. Determining that no conflict exists at row $3$, the invocation proceeds to resolve the conflict at row $6$.
@@ 4.1.1. Case with Dimension a Power of $2$
Equation 6 indicates that $U^{t}_{R}$ swaps the positions of $A^{\prime}_{0,1}$ and $A^{\prime}_{1,0}$, while $U^{t}_{L}$ performs block-wise transposition. This follows that:
@@
holds for $\ell\leq\log{d}-1$ if the following process defines the permutations in it:
@@
{±$\begin{split}&\{\pm(d(q-p)+p-q)\mid 0\leq p<q<d/2^{\ell})\}=\\ &\{\pm i\cdot(d-1)\mid 0\leq i<d/2^{\ell}\}.\end{split}$
Now, it is evident that Equation 7 with $U^{t}_{R_{i}}$ and $U^{t}_{\ell}$ defined above is a $-\ell$ ideal decomposition of $U^{t}$ for $11\leq\ell\leq\log{d}-$ 1.
@@
Let $r=(u+x)d+(v+y)-(p+x)d-(q+y)=d(u-p)+(v-q)$. Moving $A^{\prime}_{i,j}$ to the position of $A^{\prime}_{k,l}$ requires a rotation by step $-r$, while moving $A^{\prime}_{k,l}$ to the position of $A^{\prime}_{i,j}$ requires rotation by step $r$. Entries outside the blocks maintain their original positions requiring rotation by step 00. Furthermore, if either $A^{\prime}_{i,j}$ or $A^{\prime}_{k,l}$ is entirely zero, the corresponding rotation for that block can be omitted. Thus, the swapping involves at most three non-zero diagonals in its matrix representation. ∎
@@
To move $A^{\prime}[p,q]$ to the position of $A^{\prime}[u,v]$, it requires a rotation of $d(s+u)+(t+v)-d(s+p)-(t+q)=d(p-u)+(q-v)$ steps. It is evident that the number of different rotations for movements within the block is independent of the block’s position within $A$. ∎
@@ 4.2. Homomorphic Matrix Multiplication
Let $A$ and $B$ be the $d\times d$ matrices to be multiplied, and let the plaintext space be a vector space of dimension $n=d^{3}$. The method proposed by Rizomiliotis and Triakosia first appends $(d-1)$ zero matrices of size $d\times d$ to both the end of $A$ and $B$ to form $\hat{A}$ and $\hat{B}$ of size $d^{2}\times d$, where each of them can be encrypted by a ciphertext. Two permutations $U^{\gamma}$ and $U^{\xi}$ are then applied to $\hat{A}$ and $\hat{B}$ respectively. Next, column-wise replication is performed on $U^{\gamma}(\hat{A})$ and row-wise replication on $U^{\xi}(\hat{B})$. Denote the replicated results as $\tilde{A}$ and $\tilde{B}$. Finally, compute $\hat{C}=\tilde{A}\odot\tilde{B}$, and perform $\hat{C}\leftarrow\texttt{Rot}(\hat{C},d^{3}/2^{i})+\hat{C}$ for $1\leq i\leq\log{d}$. This yields the desired result $C=AB$ in the first $d\times d$ block of $\hat{C}$.
@@ 4.2.1. Decomposability of $U^{\gamma}$ and $U^{\xi}$
The rule from the previous section can be applied to this transposition. Thus, the decomposition of $U^{\gamma}$ is similar to $U^{t}$. Specifically, when $d$ is a power of two, $U^{\gamma}$ has a $-\ell$ ideal decomposition for $\ell\leq\log{d}-1$. When $d$ is set arbitrarily, $U^{\gamma}$ has an approximate $-\ell$ ideal decomposition for $\ell\leq\lfloor\log{d}\rceil-1$.
Furthermore, we can modify the decomposition of $U^{\gamma}$ to achieve constant multiplication depth. We fill the non-zero diagonals of $U^{\gamma}$ with $1$ and denote the result as $U^{\gamma}_{pad}$ (proof provided in Appendix B.2). At this point, $U^{\gamma}_{pad}(\hat{A})$ is the sum of the rotated $\hat{A}$ by steps of $0,-(d^{2}-1),\dots,-(d^{2}-1)(d-1)$ without any masking. It can be observed that $U^{\gamma}_{pad}(\hat{A})$ has the same first column as $U^{\gamma}(\hat{A})$. This implies that we can get $U^{\gamma}(\hat{A})$ by masking $U^{\gamma}_{pad}(\hat{A})$ only once. The decomposition of $U^{\gamma}_{pad}$ can be directly obtained by filling all non-zero diagonals in the permutation matrices of $U^{\gamma}$’s decomposition with $1$, yielding a constant multiplication depth of $1$.
@@ 4.2.2. Flexible Utilization of the Encoding Space
The original algorithm by Rizomiliotis and Triakosia necessitates a plaintext encoding space of $d^{3}$ to perform a $d\times d$ matrix multiplication, resulting in a fast-growing computational overhead as $d$ increases. Here, we propose a strategy to enable configurable encoding space. In essence, we split $U^{\gamma}(\hat{A})$ and $U^{\xi}(\hat{B})$ each into $d/d^{\prime}$ blocks (each size is $d^{\prime}\times d^{\prime}$ vectors), reducing the encoding space required for representing $A$ and $B$ to $d^{2}\cdot d^{\prime}$. Let $d^{\prime}$ be a divisor of $d$, $\hat{A}$ and $\hat{B}$ are both interpreted as the concatenation of $d/d^{\prime}$ blocks as follows:
@@
where $\hat{G}=\hat{A},\mathbf{g}_{i}=\mathbf{a}_{i},\mathbf{0}=\mathbf{0}_{c}$ (or $\hat{G}=\hat{B},\mathbf{g}_{i}=\mathbf{b}_{i},\mathbf{0}=\mathbf{0}_{r}$). To complete an HMM, we apply block-wise $U^{\gamma}$ and $U^{\xi}$ to each $\hat{A}^{\prime}_{i}$ and $\hat{B}^{\prime}_{i}$ respectively, yielding: $U^{\gamma^{\prime}}(\hat{A})=[U^{\gamma}(\hat{A}^{\prime}_{0})\ldots U^{\gamma}(\hat{A}^{\prime}_{d/d^{\prime}-1})]$ and $U^{\xi^{\prime}}(\hat{B})=[U^{\xi}(\hat{B}^{\prime}_{0})\ldots U^{\xi}(\hat{B}^{\prime}_{d/d^{\prime}-1})]$. This costs $2\log{d^{\prime}}$ rotations in total (Theorem 2). For each $U^{\gamma}(\hat{A}^{\prime}_{i})$ (or $U^{\xi}(\hat{B}^{\prime}_{i})$), we use masking to extract it then replicate its first colum to the entire encoding space individually, yielding the output $\tilde{A}_{i}$ (or $\tilde{B}_{i}$). This costs $2\frac{d}{d^{\prime}}\log{d}$ rotations in total. Finally, we compute $\hat{C}=\sum_{i=0}^{\frac{d}{d^{\prime}}-1}\tilde{A}_{i}\odot\tilde{B}_{i}$, followed by a summation over $\log d^{\prime}$ rotations on $\hat{C}$ to output the desired result. The total number of rotations is $3\log d^{\prime}+2\frac{d}{d^{\prime}}\log d$.
On the other hand, a ciphertext may store more matrices as the encoding space requirement decreases. For $m$ pairs of $d\times d$ matrices $A_{i}$ and $B_{i}$ ($0\leq i<m$) stored in $\hat{\mathcal{A}}=[\hat{A}_{0},\ldots,\hat{A}_{m-1}]^{T}$, the HMM between $\hat{\mathcal{A}}$ and $\hat{\mathcal{B}}$ yields $\hat{\mathcal{C}}=[\hat{C}_{0},\ldots,\hat{C}_{m-1}]^{T}$, where each $\hat{C}_{i}$ contains $C_{i}=A_{i}B_{i}$. Such a parallelizable setup gives an amortized rotation complexity of $\left(\frac{3}{m}\log d^{\prime}+2\frac{d}{d^{\prime}m}\log d\right)$. Further trade-offs are discussed in Appendix B.3 for reducing the complexity component $(\frac{d}{d^{\prime}}\log{d})$ of replications to $\left(\frac{d}{d_{0}}+\frac{d}{d^{\prime}}\log{d_{0}}\right)$ for $1\leq d_{0}\leq d$.
@@
(Jiang et al., 2018) $3d+5\sqrt{d}$ $d$ $3$ $d^{2}$ $\frac{n}{d^{2}}$
(Rizomiliotis and Triakosia, 2022) $2d+3\log d$ $1$ $2$ $d^{3}$
(Huang and Zong, 2023) $d\log{d}+d$ $d$ $2$ $d^{2}$
(Zhu et al., 2023) $4d$ $2d$ $2$ $d^{2}$
Ours $2d\left(\frac{1}{d_{0}}+\frac{\log{d_{0}}}{d^{\prime}}\right)+3\log{d^{\prime}}$ $\frac{d}{d^{\prime}}$ 2 $d^{2}d^{\prime}$ $\frac{n}{d^{2}d^{\prime}}$

### v7

@@ 3.2. Concrete Design
Based on the observation above, we design an algorithm searching a depth-1 ideal decomposition $U=U_{L}U_{R}$ for any $n\times n$ permutation $U$. The algorithm consists of three steps, as outlined below. Its pseudocode is provided in Algorithm 1.
Step 1. (Decompose Entries) This step decomposes each non-zero entry in $U$ to the $r_{c}$-th, $r_{c}-$th, or 0-th diagonals of $U_{R}$. $r_{c}$ corresponds to $r_{c}^{(1)}=a\cdot\lceil\frac{r/a}{2}\rceil$ in Definition 2, where $U$’s non-zero diagonal indices lie in [r,r][-r,r] and $a$ is the greatest common difference found. First, we create a set of key-value tables representing the non-zero diagonals of $U_{R}$:
@@
where $M_{i}$ represents the $i$-th non-zero diagonal for $i\in\{\pm r_{c},0\}$. For each table, its keys indicate the row coordinates of entries on it, and its values denote the diagonal coordinates of entries in $U$ determined by the keys. Next, for each non-zero entry $U(k,l)^{*}$ in $U$:
@@
    If $r_{c}<k<r_{c}$, add $\langle\texttt{ConvertR}(k,l,0),k\rangle$ to $M_{0}$;
@@
According to Theorem 1 and Definition 2, for the diagonal distribution of $U_{L}$ to converge to $r+r_{c},r-r_{c}]$ with $r_{c}=a\cdot\lceil\frac{r/a}{2}\rceil$, non-zero entries on the $k$-th diagonal of $U$ with $|k|>r_{c}$ can only be decomposed to the $\texttt{sign}(k)\cdot r_{c}$-th diagonal of $U_{R}$. Step 1 employs the first two conditional statements to receive these entries for decomposition. If conflicts arise between the $r_{c}$-th and $r_{c}-$th diagonals during the decomposition of these entries, it indicates that no ideal decomposition exists for $U$. Such cases are captured by the fourth conditional statement.
The remaining non-zero entries of $U$ are distributed on diagonals whose indices’ absolute value do not exceed $r_{c}$, and these entries can either be decomposed to the 0-th diagonal or the nearest diagonal among the $r_{c}$-th and $r_{c}-$th diagonals according to Theorem 1. The strategy in Step 1 is to decompose all of them to the 0-th diagonal.
After Step 1, $U_{R}$ has non-zero entries only on the 0-th and $\pm r_{c}$-th diagonals, with potential conflicts occurring only between the 0-th and $r_{c}$-th diagonals, as well as between the 0-th and $r_{c}-$th diagonals.
@@
Assigned with a specific row index $a$, the function CheckRow first checks if any conflict exists in row $a$. If there is a conflict, it attempts to resolve the conflict; otherwise, it proceeds to resolve the next conflict in $Q$. From the previous step, it is known that the conflicts occur between the 0-th diagonal and the $\pm r_{c}$-th diagonals. Therefore, for any conflicting row $a$, CheckRow can only resolve it by moving the entry $U_{R}(0,a)^{*}$ to the $r_{c}$-th (or $r_{c}-$th) diagonal. The condition for successfully moving $U_{R}(0,a)^{*}$ is that the target row $\hat{a}$ must not contain any entries on the $r_{c}$-th (or $r_{c}-$th) diagonal. If such an entry exists (with a corresponding key-value pair $\langle\hat{a},v\rangle$ in $M_{k}$ for $k=r_{c}$ or $r_{c}$), two possible cases arise:
@@
[…]
@@
[…]
Therefore, $U_{R}(0,a)^{*}$ should not be moved to row $\hat{a}$ in either of these cases. If the movement is successful, CheckRow invokes itself to resolve any potential conflict between the $r_{c}$-th (or $r_{c}-$th) diagonal and the 0-th diagonal for row $\hat{a}$. If the invocation returns 0 for irresolvable conflicts, then CheckRow has to undo the movement it employs and return 0 to the previous recursive call. If the invocation determines no conflict exists in row $\hat{a}$, it proceeds to resolve the next conflict recorded in $Q$. In this way, CheckRow operates as a depth-first search, exhaustively checking all entry movement combinations until it identifies the first conflict-free $U_{R}$.
@@
The correctness of the search algorithm in detecting whether a given permutation $U$ admits any ideal depth-1 decomposition is demonstrated throughout the description above. In the next section, the search algorithm is applied to specific permutations in homomorphic matrix operations, where we provide a visualizable demonstration of this algorithm.
Note that a given permutation may admit none, one, or multiple ideal depth-1 decompositions, and the same holds for each of its possible resulting factors $U_{L}$. Thus, a straightforward recursive invocation of our search algorithm does not guarantee an ideal decomposition of maximal depth. However, the search algorithm is capable of finding all ideal depth-1 solutions for $U$. This is achieved by slightly adapting the algorithm to treat the final resolved conflict as unresolvable every time, prompting it to backtrack through previous conflicts for alternative solutions. By further applying the search algorithm to the $U_{L}$s of all identified solutions, we can determine whether $U$ admits a depth-$2$ decomposition. This approach naturally extends to a depth-first search strategy for obtaining $U$’s maximal ideal decomposition depth.
@@ 4.1. Homomorphic Matrix Transposition
Let $ct(\mathbf{a})$ denote a ciphertext of an $n$-dimensional vector $\mathbf{a}$, where $\mathbf{a}$ is the row-ordering encoding of a $d\times d$ matrix AA ($d^{2}\leq n$). To obtain the ciphertext of the row-ordering vector of $A^{T}$, a specific homomorphic permutation (denoted as $n\times n$ matrix $U^{t}$) is applied to $ct(\mathbf{a})$. $U^{t}$ possesses 2d12d-1 nonzero diagonals defined as:
[…]
where $j\in[0,d-|i|)$. Currently, the BSGS-based HLT mentioned previously yields the best performance for $U^{t}$ with an overhead of approximately $d_{1}+2d_{2}$ rotations and keys, where $d_{1}d_{2}=d$. As noted previously, using the Benes network for decomposition would yield a rotation complexity no less than $O(4\log{d^{2}})$, which is inefficient within the range $0<d<2^{8}$ for practical HE parameter settings (Bossuat et al., 2024), not to mention the required $2(\log{d^{2}}-1)$ multiplicative depth.
4.1.1. Case with Dimension a Power of 22
For $d$ a power of $2$, $U^{t}$ can be precisely decomposed to $U^{t}_{L}U^{t}_{R}$ when we apply the depth-$1$ decomposition searching algorithm proposed in Section 3. This procedure is illustrated in Figure 3, which depicts the decomposition search of a $U^{t}$ with size $16\times 16$ ($d=4$).
The $r_{c}$ is set to $(d-1)\cdot\lceil\frac{(d-1)^{2}/(d-1)}{2}\rceil=6$ as the indices of non-zero diagonals in $U^{t}$ have a common difference of d1=3d-1=3. Step 1 of the algorithm is presented in Figure 3b. We can observe that $U^{t}$’s entries in diagonals $k\geq 6$ and $k\leq-6$ are decomposed to the $6$-th and $6-$th diagonal of $U_{R}^{t}$ respectively, while those in diagonal 6<k<6-6<k<6 are decomposed to the 0-th diagonal. This introduces conflicts in rows $6$ and $9$. Figure 3c shows how Step 2 resolves the conflicts, illustrated by the dash paths. For example, let CheckRow begin with row $9$: CheckRow first moves the conflicting entry $U_{R}(0,9)$ to the $6$-th diagonal as $M_{0}[9]>0$, and this corresponds to a destination of row $3$. CheckRow then invokes itself to check row $3$. Determining that no conflict exists at row $3$, the invocation proceeds to resolve the conflict at row $6$.
@@ 4.1.1. Case with Dimension a Power of 22
Equation 6 indicates that $U^{t}_{R}$ swaps the positions of $A^{\prime}_{0,1}$ and $A^{\prime}_{1,0}$, while $U^{t}_{L}$ performs block-wise transposition. It follows that the equation
@@
holds for $\ell\leq\lfloor\log{(d-1)}\rfloor$ if the process below defines the permutations in the equation:
@@
$\begin{split}&\{\pm(d(q-p)+p-q)\mid 0\leq p<q<d/2^{\ell})\}=\\ &\{\pm i\cdot(d-1)\mid 0\leq i<d/2^{\ell}\}.\end{split}$
Now, it is evident that Equation 7 with $U^{t}_{R_{i}}$ and $U^{t}_{\ell}$ defined above is a depth-$\ell$ ideal decomposition of $U^{t}$ for $1\leq\ell\leq\lfloor\log{(d-1)}\rfloor$. The concrete rotation complexity for evaluating a depth-$\ell$ decomposed $U^{t}$ is $2\ell+2d_{2}+d_{1}$. Specifically, the factor $U_{L_{\ell}}$ has around $2d/2^{\ell}$ non-zero diagonals symmetrically distributed on both sides of the 0-th diagonal, and is computed using HLT with $d_{1}$ BSGS inner loops and $2d_{2}$ outer loops. Other factors, $U_{R_{1}},...,U_{R_{\ell}}$, are each computed via HLT without the BSGS algorithm, contributing $2\ell$ rotations.
@@
Let r=(u+x)d+(v+y)(p+x)d(q+y)=d(up)+(vq)r=(u+x)d+(v+y)-(p+x)d-(q+y)=d(u-p)+(v-q). Moving $A^{\prime}_{i,j}$ to the position of $A^{\prime}_{k,l}$ requires a rotation by step r-r, while moving $A^{\prime}_{k,l}$ to the position of $A^{\prime}_{i,j}$ requires rotation by step $r$. Entries outside the blocks maintain their original positions requiring rotation by step 0. Furthermore, if either $A^{\prime}_{i,j}$ or $A^{\prime}_{k,l}$ is entirely zero, the corresponding rotation for that block can be omitted. Thus, the swapping involves at most three non-zero diagonals in its matrix representation. ∎
@@
To move $A^{\prime}[p,q]$ to the position of $A^{\prime}[u,v]$, it requires a rotation of d(s+u)+(t+v)d(s+p)(t+q)=d(pu)+(qv)d(s+u)+(t+v)-d(s+p)-(t+q)=d(p-u)+(q-v) steps. It is evident that the number of different rotations for movements within the block is independent of the block’s position within $A$. ∎
@@ 4.2. Homomorphic Matrix Multiplication
[…]
@@ 4.2.1. Decomposability of $U^{\gamma}$ and $U^{\xi}$
The rule from the previous section can be applied to this transposition. Thus, the decomposition of $U^{\gamma}$ is similar to $U^{t}$. Specifically, when $d$ is a power of two, $U^{\gamma}$ has a depth-$\ell$ ideal decomposition for $\ell\leq\lfloor\log{(d-1)}\rfloor$. When $d$ is set arbitrarily, $U^{\gamma}$ has an approximate depth-$\ell$ ideal decomposition for $\ell\leq\lfloor\log{d}\rceil-1$.
[…]
@@ 4.2.2. Flexible Utilization of the Encoding Space
[…]
@@
where $\hat{G}=\hat{A},\mathbf{g}_{i}=\mathbf{a}_{i},\mathbf{0}=\mathbf{0}_{c}$ (or $\hat{G}=\hat{B},\mathbf{g}_{i}=\mathbf{b}_{i},\mathbf{0}=\mathbf{0}_{r}$). To complete an HMM, we apply block-wise $U^{\gamma}$ and $U^{\xi}$ to each $\hat{A}^{\prime}_{i}$ and $\hat{B}^{\prime}_{i}$ respectively, yielding: $U^{\gamma^{\prime}}(\hat{A})=[U^{\gamma}(\hat{A}^{\prime}_{0})\ldots U^{\gamma}(\hat{A}^{\prime}_{d/d^{\prime}-1})]$ and $U^{\xi^{\prime}}(\hat{B})=[U^{\xi}(\hat{B}^{\prime}_{0})\ldots U^{\xi}(\hat{B}^{\prime}_{d/d^{\prime}-1})]$. This costs $2\log{d^{\prime}}$ rotations in total (Theorem 2). For each $U^{\gamma}(\hat{A}^{\prime}_{i})$ (or $U^{\xi}(\hat{B}^{\prime}_{i})$), we use masking to extract it then replicate its first colum to the entire encoding space individually, yielding the output $\tilde{A}_{i}$ (or $\tilde{B}_{i}$). These $2\frac{d}{d^{\prime}}$ replications cost $2\frac{d}{d^{\prime}}\log{d}$ rotations in total. Finally, we compute $\hat{C}=\sum_{i=0}^{\frac{d}{d^{\prime}}-1}\tilde{A}_{i}\odot\tilde{B}_{i}$, followed by a summation over $\log d^{\prime}$ rotations on $\hat{C}$ to output the desired result. The total number of rotations is $3\log d^{\prime}+2\frac{d}{d^{\prime}}\log d$.
In addition, a ciphertext may store more matrices as the encoding space requirement decreases. For $m$ pairs of $d\times d$ matrices $A_{i}$ and $B_{i}$ ($0\leq i<m$) stored in $\hat{\mathcal{A}}=[\hat{A}_{0},\ldots,\hat{A}_{m-1}]^{T}$, the HMM between $\hat{\mathcal{A}}$ and $\hat{\mathcal{B}}$ yields $\hat{\mathcal{C}}=[\hat{C}_{0},\ldots,\hat{C}_{m-1}]^{T}$, where each $\hat{C}_{i}$ contains $C_{i}=A_{i}B_{i}$. Such a parallelizable setup gives an amortized rotation complexity of $\left(\frac{3}{m}\log d^{\prime}+2\frac{d}{d^{\prime}m}\log d\right)$. Further trade-offs are discussed in Appendix B.3 for reducing the complexity component $(\frac{d}{d^{\prime}}\log{d})$ of replications to $\left(\frac{d}{d_{0}}+\frac{d}{d^{\prime}}\log{d_{0}}\right)$ for $1\leq d_{0}\leq d$.
@@
(Jiang et al., 2018) $5\sqrt{d} dd 33$ $d^{2}$ $\frac{n}{d^{2}}$
(Rizomiliotis and Triakosia, 2022) $2d+3\log d$ 11 22 $d^{3}$ -
(Huang and Zong, 2023) $\log{d}+d dd$ 22 $d^{2}$ -
(Zhu et al., 2023) $4d$ $2d$ 22 $d^{2}$ -
Ours $3\log{d^{\prime}}+2\frac{d}{d^{\prime}}\log{d}$ $\frac{d}{d^{\prime}}$ 2 $d^{2}d^{\prime}$ $\frac{n}{d^{2}d^{\prime}}$
@@
Figure 4. A multi-group network instance constructed for a length-$16$ permutation